Fermi National Laboratory


Accelerator Update

Wednesday April 14
The day shift began with Operations monitoring store 3399 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiment taking beam.

The Neutron Therapy Facility (NTF) began running studies at 8:10 AM. The studies ran all day.

Operations reported problems with a Pbar Debuncher RF station (DRF1) at 10:19 AM. Experts investigated and found a temperature problem with the momentum optical notch filter. This slowed stacking and eventually caused a small loss of the stack.

A number of Main Injector RF stations tripped off within a few minutes of 10:31 AM (MIRF5, 6, 8, 18). They all reset okay.

At 10:35 AM, MiniBooNE began a Horn-off run. (The experiment will run this way for several days.)

Operations reported at 10:40 AM that they were having problems with the Meson line SWICs and SEMS. An expert investigated. At 2:36 PM, a technician found a problem with a high voltage supply and found two chambers with bad wires.

MI, Recycler, and Pbar began a study period at 1 PM.

At 3:21 PM, operators found a MCenter critical device power supply (MC1D) not regulating properly.

Stacking resumed at 4:14 PM.

An MI expert began a five-hour barrier bucket study at 4:50 PM.

At 6:39 PM, a Linac RF station (LRF3) tripped off and wouldn't reset. Operators changed out a module. Pbar resumed stacking at 7:28 PM.

A power supply tech repaired MW1W at 8:29 PM.

Operators found a problem with the power amplifier tube for LRF3 at 8:51 PM. They shut off the station and called in a Linac tech who made repairs.

At 9:30 PM, TeV experts began an end-of-store study.

MTest finished using beam for the night at 9:44 PM.

A power supply tech replaced at firing module for MC1D and had the supply back in operation by 10:36 PM.

Operations terminated store 3399 at 10:38 PM and had the TeV back in shot setup by 11:33 PM.

Thursday April 15
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

At 1:58 AM, Operations established store 3401 with an initial luminosity of 50.78E30.

At 3:08 AM, the TeV suffered a multihouse quench due to a kicker prefire.

A vacuum problem in the Linac 400 MeV line halted all downstream beam. Operators found some closed valves, cycled them and had beam restored by 5:56 AM.

At 7:27 AM, the Run Coordinator decided that there would be no MI/Recycler access as previously planned. This will be rescheduled for some later date.

The experimenters at the Collider Detector at Fermilab (CDF) reported at 7:44 AM that they were having problems with their silicon detector. Repairs would take from an hour to an hour and a half.

The Pbar lithium lens tripped off at 9:38 AM. Operators had the lens reset within ten minutes, but then it tripped off again two minutes later. Operators called an expert who had the problem solved by 10:04 AM.

At 10:50 AM, the Run Coordinator decided to allow a TeV abort kicker checkout. This checkout would take about eight hours.

Both CDF and the D-zero (D0) experiments accessed their collision halls around 12:18 PM. They were both out of access by 3:15 PM.

CDF had to go back in on access at 4 PM. They came out at 5:10 PM.

Operations put the TeV through a dry squeeze at 6:05 PM.

At 6:55 PM, Operations reported they couldn't get the Tevatron Electron Lens (TEL) on. An expert replaced a screen power supply and had the TEL on within six minutes.

Operations established beam to MCenter at 8:05 PM.

MI studies ended for the night at 8:16 PM.

Operations put the TeV through another dry squeeze at 8:53 PM.

At 9:50 PM, while Operations was putting the TeV through a wet squeeze there was a kicker (AAK1) prefire fail to fire error. Operators contacted experts.

At 11:30 PM, the TeV experts reported that they had to change out the trigger source and the controls power supply for AAK1, which would take about two hours. TeV experts would be allowed to conduct studies for the rest of the night. Once repaired, the kicker would run all night and then be re-evaluated in the morning.

An MI expert began a coalescing study at 11:34 PM.

Friday April 16
The midnight shift began with TeV kicker repairs in progress, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

A Pbar Accumulator kicker module (A:IKIK) stopped firing at 1:31 AM. Operators called in an expert, but at 2:03 AM, just before he arrived, the module started firing normally.

At 1:47 AM, the kicker experts reported that they had completed their work on AAK1 and set up a scope to monitor the high voltage trigger.

TeV experts began studies at 1:54 AM.

Plans for Today and the Weekend
The TeV will continue to conduct studies until repairs begin on the AAK1 trigger electronics later this morning. The repairs are expected to take four hours. Later today, the Recycler will stash antiprotons from Pbar and Booster experts will conduct cogging studies for slip stacking operations. After all that gets completed, stacking and stores.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/16/2004   email Fermilab